Project

General

Profile

03042014 Minutes

Attending: Michael, Mat, Dominick

-Sudden dropping hits per spill observed yesterday morning at 7 am is running stably at new lower level.

-Observed lots of short stubby tracks in P2 that didn't span longer than a DCM. Evan fixed this with a timing update. Syncs were "stuck".

-Partition 3 - doesn't seem to be keeping up. Michael will open up the "valve".

-Partition 1 - sync issue at 3 pm yesterday.

Michael -
- Watch dog page shows that
NDOS Nearline blank - bug from new release - no cosmic stream. (Datalogger broken?) Louise is fixing this.
DDT streams - Michael is worried that find will break.

Joao - We need to ask about reco metric.